OpenStudy (naudiapie):

A jacket with one patch at Hot Wear Jackets costs $88.00 and each additional patch is $3.00. The same jacket at ABC Clothing costs $66.95 and each additional patch is $4.50. With a certain number of patches the jackets cost the same. Which jacket costs more with additional patches?

OpenStudy (johnweldon1993):

\[\large y_1 = 3.00x + 88.00\] \[\large y_2 = 4.50x + 66.95\] We can find the number of patches it would take to make these 2 jackets cost the same by setting these 2 equations equal to each other and solving for 'x' \[\large 3.00x + 88.00 = 4.50x + 66.95\] \[\large 1.50x = 21.05\] \[\large x = 14.03\] So after 14 patches they cost the same...anything more than that, The jacket at ABC will begin to cost more because the cost per patch is higher

OpenStudy (naudiapie):

wow that's what I got but it sounded so wrong

OpenStudy (naudiapie):

i did 88 -66.95 and 4.50 -3

OpenStudy (naudiapie):

then divided

OpenStudy (naudiapie):

then it came up with 14.03333333333 on calculator but i thought i was wrong lol

OpenStudy (johnweldon1993):

Nope that's correct :) I know, they made it so it wouldn't make sense -__- haha should be a nice number >.< but oh well...math can be a butt sometimes :P

OpenStudy (naudiapie):

Can you help me with one more please?

OpenStudy (johnweldon1993):

No >.<

OpenStudy (johnweldon1993):

Haha of course :D

OpenStudy (naudiapie):

/cries

OpenStudy (naudiapie):

awe yay okay

OpenStudy (naudiapie):

Nathan Alexander wants to purchase ears of corn for a family reunion picnic. A farmers’ market near his home offers six ears of corn for $1.25, 12 ears of corn for $2.64, or 36 ears of corn for $7.20. Which is the best buy?

OpenStudy (johnweldon1993):

Well since I know you knew what to do last time...you tell me :D Show me whatcha got! lol

OpenStudy (naudiapie):

you don't have to give me the answer , i just have trouble setting up an equation

OpenStudy (naudiapie):

7.20 - 2.64 - 1.25?

OpenStudy (johnweldon1993):

Not quite :) This one, we dont need any equations...they simply want to know which option is better...heres an example If I offer you 2 pieces of chocolate...for 2 dollars and then someone else offers you 4 pieces of chocolate for 8 dollars Who would you take?

OpenStudy (naudiapie):

2 for 2 pieces because if you wanted to get 4 pieces it would cost you 4 instead of 8

OpenStudy (johnweldon1993):

Exactly...but work backwards My 2 pieces for 2 dollars...means each only costs 1 dollar The other guys 4 pieces for 8 dollars means each piece costs 2 dollars So my offer is obviously better

OpenStudy (naudiapie):

it's just that the numbers are too big for me to do mental math, that's why I asked for an equation...

OpenStudy (johnweldon1993):

so HERE 6 ears of corn for 1.25 means each ear costs how much? 12 ears of corn for 2.64 means each ear costs how much? and 36 ears of corn for 7.20 means each ear costs how much? Then you just choose which is better :)

OpenStudy (naudiapie):

Eughh

OpenStudy (naudiapie):

My question is what do i use to figure it out though, It's hurting my brain

OpenStudy (naudiapie):

It just makes no sense because it's not telling me the max amount of ears I need to get

OpenStudy (naudiapie):

when you asked me your max was 8

OpenStudy (johnweldon1993):

You use that type of thinking...no equations this time around :) Not trying to hurt your brain hun >.< So here 6 ears of corn...for 1.25 means each ear of corn cost \(\large \frac{1.25}{6} = 0.208\) 12 ears of corn...for 2.64 means each ear of corn cost \(\large \frac{2.64}{12} = 0.22\) and 36 ears of corn...for 7.20 means each ear costs \(\large \frac{7.20}{36} = 0.20\) So in THIS case...buying the 36 ears at 7.20 is BEST because you are paying less per ear of corn :)

OpenStudy (naudiapie):

wouldn't that be an equation though, you are dividing right?

OpenStudy (johnweldon1993):

Haha well I guesssssss thats an equation

OpenStudy (naudiapie):

You divided 7.20 by 36 , that was my question all along, what do i use to to figure out the differences and it was division, that was literally all i needed to know

OpenStudy (naudiapie):

now i feel stupid

OpenStudy (johnweldon1993):

>.< I totally overlooked that then lol....dont feel stupid at all that was my fault!

OpenStudy (naudiapie):

xD <333

OpenStudy (johnweldon1993):

Sheesh, I guess I just like the numbers so much XD who needs equations :P

OpenStudy (naudiapie):

shall we do one more just for safe measures, this time I will try to do the work but you can help me check it

OpenStudy (johnweldon1993):

Haha oh for safe measures indeed ;p sounds good!

OpenStudy (naudiapie):

okay so this is the question Mel Garrett needs four new tires for his car. He can buy one tire for $87, two tires for $174, three tires for $261, or four tires for $348. Which is the best buy? All tires are the same price. two for $174 one for $87 three for $261

OpenStudy (naudiapie):

The answer is all tires are the same because (87)/(1)= 87 (174)/(2)=87 (348)/(4)=87

OpenStudy (johnweldon1993):

\[\Huge \color \red{\checkmark}\] *Throws mini party!*

OpenStudy (naudiapie):

awee c:

OpenStudy (johnweldon1993):

See you get a balloon too! Haha 

OpenStudy (naudiapie):

omg lmao, that's cute , did you draw it ? :D

OpenStudy (johnweldon1993):

of course :D haha Just for you /.\

OpenStudy (naudiapie):

o/////o /fullbodyblush

OpenStudy (johnweldon1993):

Hahahaha nooo don't do that!!!! no blushing /.\ *hides*

OpenStudy (naudiapie):

lmao i can't help it... I'm naudia btw

OpenStudy (naudiapie):

you should pm me

OpenStudy (johnweldon1993):

Maybe I already did /.\